# Penetrance Estimation in Family-based Studies with the penetrance R package

### v2

@@ Abstract
Reliable methods for penetrance (age-specific risk among those who carry a genetic variant) estimation are critical to improving clinical decision making and risk assessment for hereditary syndromes. We introduce penetrance, an open-source R package, to estimate age-specific penetrance using family-history pedigree data. The package employs a Bayesian estimation approach, allowing for the incorporation of prior knowledge through the specification of priors for the parameters of the carrier distribution. It also includes options to impute missing ages during the estimation process, addressing incomplete age information which is not uncommon in pedigree datasets. Our software provides a flexible and user-friendly tool for researchers to estimate penetrance in complex family-based studies, facilitating improved genetic risk assessment in hereditary syndromes. The penetrance R package is freely available on CRAN, with accompanying documentation and examples.
@@ 1 Introduction
In this paper, a “PGV” refers specifically to a germline allele known or strongly suspected to be pathogenic. Risk counseling can be performed at the level of the specific PGV, but more often the interest lies in estimating risk for groups of PGVs that occur in the same gene, gene region, or group of genes. For example, one may be interested in the risk associated with the I307K mutation in the APC gene, any PGV in the APC gene, or any PGV in a group of genes comprising Lynch syndrome. Here, we define a ”carrier” as an individual who has a PGV within a specific set of interest. Thus our model can be applied equally to the estimation of PGV-specific, gene-specific or syndrome-specific penetrance.
A penetrance function is the proportion of individuals who develop a condition by a certain age among carriers. Because of the large number of genes probed by genetic panel testing, the relative rarity of PGVs in many of them, and the importance of investigating risk among specific ethnic subgroups, as well as penetrance modifiers, penetrance estimation is emerging as a critical bottleneck in implementing precision prevention for individuals at increased familial risk.
Penetrance estimation has been extensively studied in the past for a variety of study designs (Thomas, 2004). Family-based studies usually center around a proband who is typically the initial point of contact with the healthcare provider and is the individual reporting their family history. The proband is often the only genotyped individual in the family and may or may not report additional genotyping information for known relatives. Different methods for penetrance estimation have also been previously developed for family-based studies (Chen et al., 2009; Shin et al., 2020). Classical segregation analysis, for example, has been widely used to infer inheritance patterns and estimate penetrance, but it often requires large, well-characterized pedigrees and can yield imprecise estimates when sample sizes are small (Elston and Stewart, 1971; Claus et al., 1991).
@@
We introduce an approach and the associated open-source R software package, called penetrance, to carry out age-specific and sex-specific penetrance estimation for complex family-based studies where carrier status is not observed for all family members. The implementation is based on Bayesian estimation methods, which allows for the incorporation of prior knowledge from existing studies and provides a flexible and easy-to-use method for penetrance estimation. The penetrance function for carriers is modeled using a modified Weibull distribution.
While this approach has broad applications across many heritable conditions, our primary motivation comes from cancer genetics, where early detection and prevention strategies have shown significant impact on patient outcomes (Rebbeck et al., 2018, 2009; Finch et al., 2006). In the following we introduce the software and provide the user with an illustration of our penetrance estimation approach and package, applying it to estimate colorectal cancer penetrance among MLH1 carriers using a simulation study motivated by real data of 130 probands with PGVs in the MLH1 gene, reporting family history on 4,604 individuals (including probands).
@@ 2.1 Model
We consider a pedigree-based framework for analyzing disease risk. While some sets of PGVs may increase risk for multiple diseases, our framework focuses exclusively on analyzing the age-specific penetrance for a single disease at a time. In our model, families are indexed by $i=1,\dots,I$, with members indexed by $j=0,\dots,J_{i}$ ($j=0$ denotes the proband). For family $i$, disease status are represented by $H_{i}=(H_{i0},\dots,H_{iJ_{i}})$, where each $H_{ij}=(d_{ij},y_{ij})$ denotes individual $j$’s binary diagnosis status ($d_{ij}$) and age at diagnosis or censoring ($y_{ij}$) for a single disease.
An individual’s carrier status is denoted by $G_{ij}$, which takes a binary value to indicate whether individual $i$ is a carrier ($G_{ij}=1$) or not ($G_{ij}=0$). We assume here that homozygous carriers of the PGV are not viable. The model accommodates missing carrier status data, with $G_{i,\text{obs}}$ and $G_{i,\text{mis}}$ denoting observed and missing carrier status respectively, such that $G_{i}=(G_{i,\text{obs}},G_{i,\text{mis}})$. We derive the family-level likelihood:
$P(H_{i}\mid G_{i,\text{obs}})=\sum_{G_{i,\text{mis}}}P(H_{i}\mid G_{i})\,P(G_{i,\text{mis}}\mid G_{i,\text{obs}}).$
@@
Our methodology is grounded in fundamental genetic and statistical assumptions. The genetic model relies on critical population-level parameters, including the carrier prevalence in the general population and the baseline age-specific probabilities of developing the disease for noncarriers. The prevalence would be approximately two times the allele frequency.
We assume conditional independence of individual disease status given the complete set of carrier status for the pedigree, and assume equivalent disease status among distinct pathogenic variants within the same gene. The rare disease syndromes are modeled with dominant inheritance patterns and focus on the first occurrence of the disease in an individual. In particular, cancer syndromes mostly exhibit dominant inheritance patterns (Shin et al., 2020; King et al., 2003; Lynch and de la Chapelle, 1999). These assumptions enable the factorization of the likelihood into individual-specific components:
$P(H_{i}\mid G_{i})=\prod_{j=0}^{J_{i}}P(H_{ij}\mid g_{ij}).$
We implement the likelihood computation using the Elston-Stewart peeling algorithm through the clipp package, facilitating efficient calculation of carrier probabilities in complex pedigree structures (Golicher, 2023; Elston and Stewart, 1971).
2.2 Weibull Parameterization
We choose a Weibull distribution for our parametric estimation approach, extending it with threshold and asymptote parameters to account for minimum onset age and lifetime disease probability. The standard Weibull distribution with scale ($\alpha>0$) and shape ($\beta>0$) parameters is modified by introducing a threshold parameter ($\delta>0$), which shifts the distribution along the age axis, and an asymptote parameter ($10<\gamma<$ 1), representing the lifetime probability of developing the disease. The model’s parameters $\theta=(\alpha,\beta,\gamma,\delta)$ define the cumulative distribution function (CDF) that represents disease onset probability over time. The model can incorporate prior information through informative prior distributions, which can be specified using more interpretable conditional quartiles (see Appendix A).
2.3 Bayesian Estimation Approach
We employ a parametric Bayesian approach to find the posterior distribution of the parameters of the specified Weibull distribution, $\theta=(\alpha,\beta,\gamma,\delta)$. The direct calculation of this posterior is intractable due to the complexity of the likelihood function, which involves modeling dependencies within families and integrating over unobserved variables. Therefore, we utilize MCMC methods to approximate the posterior distribution. MCMC methods generate samples from a target distribution by constructing a Markov chain whose stationary distribution is the desired posterior distribution $P(\theta\mid H)$ (Metropolis et al., 1953; Gilks et al., 1995). Our approach uses a fine-tuned adaptive MCMC algorithm that ensures adequate exploration of the parameter space.
3 Package
@@
The methods discussed above are implemented in an easy-to-use R software package, available on CRAN. The package workflow includes three main parts: (1) the user input, including family data in the form of pedigrees, and the specifications of the options for the penetrance estimation, (2) the estimation of the posterior distribution using the MCMC approach, and (3) the output of the results in the form of the samples from the approximated posterior distribution, i.e. the estimated penetrance function. Figure 1 shows the workflow.
@@
Table 1: List of main functions in penetrance.
Name Category Description
makePriors Pre-processing Creates prior distributions based on the user inputs. Depending on the combination of user inputs, the default setting for the prior distributions are modified.
mhChain Algorithm Estimates the penetrance probability using the MCMC approach. The inputs include the formatted family pedigree data, options for the construction of the MCMC chains, and the output of makePriors.
mhLoglikelihood Algorithm Evaluates the the log-likelihood for the given data and model parameters using the clipp package for the likelihood calculation.
penetrance Main Function Runs the main function. The inputs include the formatted family pedigree data, options for the construction of the MCMC chains, and some options for the output. The default output is a list of the posterior samples for each of the parameters from each of the independently run chains.
3.1 User Input
3.1.1 Pedigree Input
@@
The main input is the family pedigree data reported by the proband. The pedigree file is an R data frame object. Each family tree is represented by a single data frame that includes the following columns: PedigreeID, ID, Sex, MotherID, FatherID, isProband, CurAge, isAff, Age, and Geno. Each row in the data frame corresponds to a unique family member, identified by a unique ID within the family and a unique PedigreeID. Connections between family members are established through the MotherID and FatherID columns, which reference the IDs of each individual’s mother and father, if this information is available. In each family tree data frame, one individual is designated as the proband indicated by the isProband column with a value of 1. In the type of family-based penetrance studies considered here, families are typically ascertained through probands who are confirmed carriers. The column CurAge indicates the censoring age, which is the current age if the person is alive or the age at death if the person is deceased, with allowed ages ranging from 1 to $max\_age$. Missing ages are allowed and can be coded as NA or left empty. The isAff column indicates the disease affection status, with 1 for diagnosed individuals, 0 for unaffected, and NA for unknown. The Age column indicates the age of disease diagnosis, encoded as NA if the individual was not diagnosed or if the age was unknown, with allowed ages ranging from 1 to $max\_age$. The Geno column includes germline testing results, i.e. the carrier status, with positive results coded 1, negative results coded 0, and unknown results as NA or left empty. Every family tree must include a PedigreeID, a numeric value representing the unique identifier for each family, ensuring that there are no duplicate entries. The Sex field must be coded as 0 for females or 1 for males, with unknown sex coded as NA. More details on the structure of the pedigree required for penetrance is available in the Fam3Pro (formerly PanelPRO) documentation where a similar structure is used (Lee et al., 2021).
3.1.2 Prior Specification
To run the MCMC algorithm described in the methods section, the prior distributions for $\theta=(\alpha,\beta,\gamma,\delta)$ need to be specified. The package provides the user with a flexible approach to prior specification, balancing customization with an easy-to-use workflow. In addition to providing default prior distributions (setting 1), the package allows users to customize the priors directly (setting 2) or based on penetrance estimates from the literature (setting 3). The following settings for the prior distribution specification are available:
  1. 1.
    Default: The default setting employs pre-specified, uninformative priors. The default setting for the priors can be seen in Table 2.
  2. 2.
    Custom Parameters: The user has the option to tune the parameters of the default setting directly by overwriting the parameters in the $prior\_params\_default$ object (see Figure 2).
  3. 3.
    Prior Elicitation from Existing Studies: The package provides additional functionality to facilitate in the prior elicitation based on existing penetrance studies. For more details on the way options to input information from published studies, see Appendix A.
(a) Default prior parameters (Setting 1)
(b) Example configuration with user-modified prior parameters (Setting 2)
Figure 2: Prior Specification.
Model Parameter Default Prior Distribution
First Quartile $Q_{25}$ Scaled Beta(6,3)
Median $Q_{50}$ Scaled Beta(2,2)
Asymptote $\gamma$ Scaled Beta(1,1)
Threshold $\delta$ Uniform(5,30)
Table 2: Default prior distributions implemented in penetrance.
Based on the user inputs, penetrance utilizes the $makePriors$ function to generate priors for each of the four Weibull distribution parameters for both males and females. The function applies the same prior distributions to both sexes, using four unique priors for all eight parameters.
3.1.3 Estimation Model Inputs
@@
The estimation procedure follows the approach outlined in the methods section. In addition to the pedigree data ($data$) and the specified priors ($prior\_params$) the user inputs a carrier prevalence ($prev$) and a baseline age-specific probability of developing the disease ($baseline\_data$). Assuming carriers are rare, the population age-specific probability of developing the disease is a good approximation for the noncarrier penetrance. Such population-level risk data usually comes from a population registry such as the Surveillance, Epidemiology, and End Results (SEER) Program for cancer in the United States (Howlader et al., 2020). Otherwise, the user may choose to calculate the noncarrier penetrance using the calculateNCPen function by setting BaselineNC = TRUE. This enables the calculation of a noncarrier penetrance as part of the estimation process, using the Weibull parameters from the current iteration. Per default, the estimation is sex-specific, estimating two distinct set of parameters for females and males. However, the user has the option to run the estimation without considering sex, by setting sex_specific = FALSE. The user can provide additional inputs to the function at run-time. An overview of the options the user can specify in the main function call can be found in Table 3.
@@
Table 3: List of model options for penetrance (excluding prior and output options).
@@
3.2 Estimation Algorithm
3.2.1 Preprocessing of Pedigree Data
The estimation requires preparation of the pedigree data, particularly handling missing age information and adjusting the data format for compatibility with the clipp package. Firstly, checks are made to validate the input data and parameters. The $transformDF$ function takes the input data in the pedigree format and transforms it for the use with penetrance and clipp. If the remove_proband parameter is set to TRUE, it removes the disease status of the probands from the data. This approach serves as a strategy for ascertainment correction, though it comes at the cost of excluding probands’ information on disease status from the sample. The appropriateness of this strategy depends on the study design and the specific mechanisms used to ascertain probands into the cohort.
3.2.2 Age Imputation
3.2.3 Algorithm Initialization
The initial set of parameters, denoted $\theta_{0}$, is drawn from the empirical distribution of the provided data, stratified by sex and disease affection status. Specifically, we calculate initial thresholds, medians, and first quartiles for both sexes based on the ages of affected individuals, ensuring that these values fall within the specified prior bounds. Asymptotes for male and female are randomly drawn between the cumulative lifetime risk of the baseline (given by the input parameter $baseline\_data$) and 1.
3.2.4 Sampling of Proposals
In each iteration of the MCMC algorithm, a new proposal is generated for the parameter vector. After generating the proposal, it is essential to ensure that each parameter falls within its respective bounds. These bounds are set to maintain realistic and meaningful parameter values based on biological plausibility (see Appendix B). If any parameter falls outside these bounds, the proposal is automatically rejected, ensuring that the algorithm only considers valid parameter values for further evaluation.
3.2.5 Likelihood Calculation
Based on the proposed parameters of the Weibull distribution, penet.fn calculates the age-specific penetrance for every individual at each iteration of the algorithm. It assumes a genetic model with a single biallelic locus. Age-specific penetrance probabilities are adjusted based on carrier status and affection status of the individual. As mentioned above, we assume that homozygous carriers are not viable.
In every iteration of the algorithm, the (log-)likelihood for the families in the pedigree file is calculated. The likelihood evaluation for a given family and penetrance matrix is performed by the $pedigree\_loglikelihood$ function from the clipp R package.
3.3 Output
The penetrance function returns a comprehensive set of outputs from the MCMC estimation process. Each chain generates its own posterior samples, and these samples collectively represent draws from the target posterior distribution. After running penetrance, the outputs consist of:
  • The posterior samples for the first quartile, median, asymptote, and threshold parameters.
  • The combined posterior samples from all chains.
  • Rejection rates for each of the chains (optionally included; see Table 4).
  • The input pedigree data for reference.
  • Summary statistics for the posterior distributions (optionally included; see Table 4).
  • Density plots of the posterior distributions for each parameter (optionally included; see Table 4).
  • Trace plots of the MCMC chains for each parameter (optionally included; see Table 4).
  • Autocorrelation function (ACF) plots for the posterior samples (optionally included; see Table 4).
  • Log-likelihood plots over iterations (optionally included; see Table 4).
  • Penetrance plots showing the resulting cumulative penetrance curves with credible intervals (optionally included; see Table 4).
  • Penetrance plots showing the probability density function (PDF) curves with credible intervals (optionally included; see Table 4).
Table 4: List of output options.
Option Name Value Default Value Definition
summary_stats TRUE/FALSE TRUE Standard summary statistics computed based on the posterior samples for each parameter.
rejection_rates TRUE/FALSE TRUE Rejection rates for each of the chains.
density_plots TRUE/FALSE TRUE Density plots showing the posterior distribution for each parameter.
plot_trace TRUE/FALSE TRUE Trace plots showing the MCMC sampling over iterations for each parameter.
plot_acf TRUE/FALSE TRUE Autocorrelation function (ACF) plots for the posterior samples of each parameter.
plot_loglikelihood TRUE/FALSE TRUE Plots of the log-likelihood values over MCMC iterations.
penetrance_plot TRUE/FALSE TRUE Penetrance plots showing the cumulative penetrance curves with credible intervals based on the posterior samples.
penetrance_plot_pdf TRUE/FALSE TRUE Penetrance plots showing the probability density function (PDF) curves with credible intervals based on the posterior samples.
probCI Float (0 to 1) 0.95 Probability level for credible intervals in penetrance plots.
4 Illustration
4.1 Data
We simulated 130 probands with PGVs in the MLH1 gene, along with a total of 4,604 individuals including probands (average family size= 35.4) with 428 CRC diagnoses (203 females and 225 males). All probands and relatives had complete age information; there were no missing age values, both for age of diagnosis for affected individuals and censoring age for unaffected individuals. All probands were assumed to be carriers, and all relatives were assumed to have unknown carrier status. The dataset, code and resulting output of this illustration can be found in the vignette of the R package.
Figure 3: Exemplary code for running penetrance for simulated data for MLH1 and CRC.
4.2 Results
We run the sex-specific estimation procedure using the default settings of the penetrance package, which include non-informative priors for the Weibull parameters, a single MCMC chain, and default proposal variances (see Figure 3). The baseline risk estimates, $baseline\_data\_default$, which approximate the noncarrier risk, are CRC incidence rates from 8 registries for the US population from 1975 to 2020 as reported by SEER (Howlader et al., 2020). The user could specify alternative priors on the parameters of the Weibull distribution as discussed in the section on prior specification. Since the simulated data was generated with MLH1 carriers, we specify the carrier prevalence for MLH1 in the function call ($prev=prevMLH1$).
We generate a chain of 20,000 posterior samples and discard the first 2,000 samples as burn-in ($burn\_in=0.1$) in each plot. We do not perform any age imputation ($age\_imputation=FALSE$ and do not attempt any ascertainment correction by removing the proband ($remove\_proband=FALSE)$. Figure 4 shows the estimated absolute and cumulative penetrance curves for females and males. The shaded areas represent the 95% credible intervals derived from the posterior samples. The penetrance function returns the outputs discussed in the package section and an R object that includes the posterior samples which can be used for subsequent analysis.
Figure 4: Age-specific absolute risk (A) and cumulative risk (B) for CRC and MLH1 for females and males with default prior parameters for simulated data.
5 Discussion
Estimating disease-specific penetrance from complex family-based studies with a limited number of families poses significant challenges. However, accurate penetrance estimation remains important as the number of studies with genotyped family members continues to grow. Penetrance estimates are a critical input to risk stratification for early detection and other decision support aspects of managing inherited susceptibility of diseases including cancer.
In this work we developed an easy-to-use, well documented, open-source R package for quantitative penetrance assessment. The software offers a user-friendly approach for estimating age-specific penetrance, particularly suitable for complex family-based studies. By using Bayesian methods, the user can integrate existing information on penetrance curves through the specification of hyperparameters on the priors of the Weibull distribution parameters. We also provide uninformative default settings that allow a user to get started with minimal input while offering an extensive list of options to customize the penetrance estimation as desired.
The package includes checks on the input data and the option to automatically impute missing age data, addressing a common challenge in family-based studies. These studies often rely on family history data, typically provided by a single participant, usually the proband, who recounts the medical history of the entire family. The run-time of the algorithm increases with the age imputation. The higher the frequency of the imputation (set with the parameter $imp\_interval$), the longer the run-time. Without the age imputation the run-time scales linearly with the number of iterations and the number of families (and family size). With small numbers of families and small family sizes, the statistical reliability of the estimates decreases and credible intervals tend to be wider. Users should exercise caution when interpreting results from small samples, and it may be advisable to increase the number of iterations to better characterize the uncertainty.
Our current software implementation uses a Weibull parameterization which assumes unimodal penetrance curves, which may not fully capture more complex, multi-modal age-specific patterns. The software currently supports the joint estimation of two sex-specific penetrance functions with a single set of Weibull parameter priors, though this framework could easily be extended to accommodate different priors for each sex as well as additional stratification dimensions. Our model is limited to evaluating a single disease, such as a specific cancer, associated with a single carrier status and does not incorporate competing risks. Individuals may, for instance, develop multiple primary cancers throughout their lifetime and are subject to various other causes of death. Future work should extend this approach to richer datasets, modeling multiple cancer types simultaneously to better assess lifetime cancer risks. Furthermore our model currently assumes that homozygous carriers are not viable, which is a reasonable assumption for many rare syndromes.
In conclusion, our open-source R package represents a significant advancement in penetrance estimation for complex family-based studies. By combining robust statistical methodology with user-friendly implementation, we have created a valuable tool for genetic epidemiology research. While acknowledging current limitations, our software provides a foundation for more sophisticated penetrance estimation approaches, ultimately contributing to improved risk assessment and clinical decision support in inherited disease management. Our initial illustration in cancer genetics demonstrates the package’s utility, and we anticipate its application across a broad spectrum of hereditary conditions.
6 Acknowledgments
We gratefully acknowledge support from the National Cancer Institute at the National Institutes of Health grants R01CA242218. We thank Dr. Chris Amos for very helpful comments.
7 Data Availability
This manuscript introduces penetrance, an innovative penetrance estimation tool. The package is available on CRAN and the source code can be accessed at https://github.com/nicokubi/penetrance
by N. Kubista, D. Braun, G. Parmigiani
Appendix A Prior Elicitation based on existing studies

### v3

# The penetrance R package for Estimation of Age Specific Risk in Family-based Studies

@@ Abstract
Reliable tools and software for penetrance (age-specific risk among those who carry a genetic variant) estimation are critical to improving clinical decision making and risk assessment for hereditary syndromes. We introduce penetrance, an open-source R package available on CRAN, to estimate age-specific penetrance using family-history pedigree data. The package employs a Bayesian estimation approach, allowing for the incorporation of prior knowledge through the specification of priors for the parameters of the carrier distribution. It also includes options to impute missing ages during the estimation process, addressing incomplete age information which is not uncommon in pedigree datasets. Our open-source software provides a flexible and user-friendly tool for researchers to estimate penetrance in complex family-based studies, facilitating improved genetic risk assessment in hereditary syndromes.
@@ 1 Introduction
A penetrance function quantifies the proportion of carriers who develop a condition by a specific age. Penetrance estimation has been extensively studied in the past for a variety of study designs (Thomas, 2004). Family-based studies usually center around a proband who is typically the initial point of contact with the healthcare provider and is the individual reporting their family history. The proband is often the only genotyped individual in the family and may or may not report additional genotyping information for known relatives. Different methods for penetrance estimation have also been previously developed for family-based studies (Chen et al., 2009; Shin et al., 2020). Classical segregation analysis, for example, has been widely used to infer inheritance patterns and estimate penetrance, but it often requires large, well-characterized pedigrees and can yield imprecise estimates when sample sizes are small (Elston and Stewart, 1971; Claus et al., 1991).
@@
We introduce penetrance, an open-source R software package, to carry out age-specific and sex-specific penetrance estimation for complex family-based studies where carrier status is not observed for all family members. The implementation is based on Bayesian estimation methods, which allows for the incorporation of prior knowledge from existing studies and provides a flexible and easy-to-use method for penetrance estimation.
Our software provides a flexible and intuitive tool for researchers to estimate penetrance in complex family-based studies, facilitating improved genetic risk assessment in hereditary syndromes. While applicable across various heritable conditions, our primary motivation stems from cancer genetics, where early detection and prevention strategies significantly impact patient outcomes (Rebbeck et al., 2018, 2009; Finch et al., 2006).
@@ 2.1 Model
Our package implements a pedigree-based Bayesian framework for estimating age-specific penetrance in family-based studies. This approach focuses on modeling disease risk for a single disease at a time, though our methods could be extended to multiple diseases in future work.
@@
The statistical model integrates critical population-level parameters, including carrier prevalence in the general population and baseline age-specific disease probabilities for non-carriers.
For penetrance modeling, we employ a modified Weibull distribution with parameters $\theta=(\alpha,\beta,\gamma,\delta)$, where scale ($\alpha>0$) and shape ($\beta>0$) are standard Weibull parameters, threshold ($\delta>0$) accounts for minimum onset age, and asymptote ($10<\gamma<$ 1) represents lifetime disease probability. This parameterization creates a flexible cumulative distribution function representing disease onset probability over time. We use a parametric Bayesian approach to estimate the posterior distribution of $\theta$. Since direct calculation is intractable due to the complexity of the likelihood function and unobserved variables, we employ a fine-tuned adaptive Markov Chain Monte Carlo (MCMC) method (Metropolis et al., 1953; Gilks et al., 1995) to approximate the posterior distribution.
2.2 Package Usage
@@
The methods discussed above are implemented in an easy-to-use R software package, available on CRAN. The package workflow includes three main parts: (1) the user input, including family data in the form of pedigrees, and the specifications of the options for the penetrance estimation, (2) the estimation of the posterior distribution using the MCMC approach, and (3) the output of the results in the form of the samples from the approximated posterior distribution, i.e. the estimated penetrance function. Figure 1 shows the package workflow.
@@
The main input is the family pedigree data reported by the proband. The pedigree file is an R data frame object. Each family tree is represented by a single data frame that includes the following columns: PedigreeID, ID, Sex, MotherID, FatherID, isProband, CurAge, isAff, Age, and Geno. Each row represents a unique family member, with family relationships established through parent ID references. The isProband column (value 1) designates the individual reporting family history. Disease status is indicated through isAff (1=diagnosed, 0=unaffected, NA=unknown) and Age (age at diagnosis or NA). Carrier status is recorded in Geno (1=carrier, 0=non-carrier, NA=unknown). The column CurAge indicates the censoring age, which is the current age if the person is alive or the age at death if the person is deceased. Missing ages are allowed and can be coded as NA or left empty. More details on the structure of the pedigree required for penetrance is available in the Fam3Pro (formerly PanelPRO) documentation where a similar structure is used (Lee et al., 2021).
To run the MCMC algorithm, the prior distributions for $\theta=(\alpha,\beta,\gamma,\delta)$ must be specified. The package provides three flexible approaches to prior specification, balancing customization with ease of use. First, the package implements reasonable default uninformative priors (First Quartile: Scaled Beta(6,3), Median: Scaled Beta(2,2), Asymptote: Scaled Beta(1,1), Threshold: Uniform(5,30)), suitable for exploratory analysis when limited prior knowledge on the shape of the penetrance curves is available. Second, users can manually adjust prior parameters by modifying the prior_params_default object, which is appropriate when researchers have specific beliefs about parameter distributions that do not align with default settings. Third, the package facilitates incorporation of published penetrance estimates through several supported formats: relative risk estimates (e.g., odds ratios) that inform the asymptote parameter, detailed age-specific risk data such as information extracted from Kaplan-Meier curves, or high-level age distribution data combined with overall study size. This literature-based approach is particularly valuable when analyzing rare variants, as it leverages existing knowledge to improve estimation precision. Supplementary Materials B provides comprehensive details on the prior elicitation process and approach used to translate published data into prior distributions.
@@
The estimation procedure follows the approach outlined in the methods section. In addition to the pedigree data ($data$) and the specified priors ($prior\_params$) the user inputs a carrier prevalence ($prev$) and a baseline age-specific probability of developing the disease ($baseline\_data$). Assuming carriers are rare, the population age-specific probability of developing the disease is a good approximation for the noncarrier penetrance. Such population-level risk data usually comes from a population registry such as the Surveillance, Epidemiology, and End Results (SEER) Program for cancer in the United States (Howlader et al., 2020). Per default, the estimation is sex-specific, estimating two distinct set of parameters for females and males. However, the user has the option to run the estimation without considering sex, by setting sex_specific = FALSE. The user can provide additional inputs to the function at run-time. An overview of the options the user can specify in the main function call can be found in the Supplementary Table A1.
The estimation requires preparation of the pedigree data, particularly handling missing age information and adjusting the data format. A penetrance function quantifies the proportion of carriers who develop a condition by a specific age. Details on how the package handles missing ages are provided in supplementary materials.
For the estimation procedure, the initial set of parameters, denoted $\theta_{0}$, is drawn from the empirical distribution of the provided data, stratified by sex and disease affection status. In each iteration of the MCMC algorithm, a new proposal is generated for the parameter vector. After generating the proposal, it is essential to ensure that each parameter falls within its respective bounds. These bounds are set to maintain realistic and meaningful parameter values based on biological plausibility (see Supplementary Materials C). If any parameter falls outside these bounds, the proposal is automatically rejected, ensuring that the algorithm only considers valid parameter values for further evaluation.
Based on the proposed parameters of the Weibull distribution, penet.fn calculates the age-specific penetrance for every individual at each iteration of the algorithm. The clipp package is used to efficiently compute the likelihoods using the Elston-Stewart algorithm (Golicher, 2023). It assumes a genetic model with a single biallelic locus. Age-specific penetrance probabilities are adjusted based on carrier status and affection status of the individual.
After running for the specified number of iterations, the penetrance function returns a comprehensive set of outputs from the MCMC estimation process. Each chain generates its own posterior samples, and these samples collectively represent draws from the target posterior distribution.
3 Illustrative Example
We simulated 130 probands with PGVs in the MLH1 gene, along with a total of 4,604 individuals including probands (average family size= 35.4) with 428 CRC diagnoses (203 females and 225 males). The dataset, code and resulting output of this illustration can be found in the vignette of the R package.
We generate a chain of 20,000 posterior samples and discard the first 2,000 samples as burn-in ($burn\_in=0.1$) in each plot. We do not perform any age imputation ($age\_imputation=FALSE$ and do not attempt any ascertainment correction by removing the proband ($remove\_proband=FALSE)$.
The penetrance function returns the outputs discussed in the methods section and an R object that includes the posterior samples which can be used for subsequent analysis.
4 Conclusion
The penetrance open-source R package is a specialized tool for estimating age-specific disease risk in carriers of pathogenic genetic variants using family-based data. It addresses the critical need for reliable penetrance estimates in genetic counseling and risk assessment for hereditary conditions. Using a Bayesian framework with a modified Weibull distribution, the package allows researchers to incorporate prior knowledge from published studies while providing sensible defaults for those without prior information.
Key features of the package include flexible prior specification options, sex-specific penetrance estimation, and automated handling of missing age data through an optional imputation algorithm. The software implements the Elston-Stewart peeling algorithm through the clipp package for efficient likelihood calculation in complex pedigree structures. Comprehensive visualization options help researchers interpret results easily.
The current implementation focuses on single-disease, single-carrier-status models with sex as the primary stratification variable, and assumes a unimodal Weibull distribution for age-specific penetrance patterns. While these design choices accommodate many hereditary syndromes, future extensions could incorporate competing risks, multiple disease outcomes, and more flexible penetrance distributions.
By combining statistical rigor with user-friendly implementation, penetrance provides genetic epidemiologists with an accessible tool for estimating disease risk from family-based studies. The package was developed in the context of cancer genetics research but is applicable across many hereditary conditions where accurate risk assessment is essential for clinical decision-making. The software is freely available on CRAN with comprehensive documentation and examples.
5 Acknowledgments
We gratefully acknowledge support from the National Cancer Institute at the National Institutes of Health grants R01CA242218. We thank Dr. Chris Amos for providing very helpful comments.
6 Data Availability
This manuscript introduces penetrance, an innovative penetrance estimation tool. The package is available on CRAN and the source code can be accessed at https://github.com/nicokubi/penetrance.
by N. Kubista, D. Braun, G. Parmigiani
Appendix A penenterance package options
@@ 6 Data Availability
Table A1: List of model options for penetrance (excluding prior and output options).
@@
Appendix B Prior Elicitation based on existing studies